# Scaling Relations of Spectrum Form Factor and Krylov Complexity at Finite Temperature

### v1

@@ Abstract
Unveiling the scaling relations of physical quantities is an important and fundamental topic in the study of classical and quantum physics Stanley (1999); Huckestein (1995); Osterloh et al. (2002); García-García and Wang (2008); Serra et al. (1998); Pal and Huse (2010); Fisher and Berker (1982). In the context of equilibrium systems, significant progress has been made in this direction, leading to a deeper understanding of critical phenomena, phase transitions, and universal behaviors Hung et al. (2011); Puebla et al. (2019); Chen et al. (2019); Chauhan et al. (2022). The investigation of scaling relations in numerous systems are still attracting a growing interests, especially in the context of non-equilibrium systems and quantum dynamics of chaotic systems Sierant et al. (2020); Šuntajs et al. (2020).
@@
Typically, the characteristics of quantum chaos manifest across the entire spectrum. To quantify the chaos strength, the statistical properties are conventionally studied. For example, the Wigner-Dyson distribution describes the statistical properties of a quantum chaotic system at infinite temperature, where a level repulsion presents in the energy spectrum Guhr et al. (1998); Mehta (1991); Haake (2010). The statistical properties of the spectrum provide information solely about the nearest level spacing within the spectrum. In contrast, the SFF measures energy correlations between any two states, which allows it to go beyond the constraints of spectrum statistics and capture richer information. SFF has found widespread use in diverse contexts, including random matrix theory (RMT) Guhr et al. (1998); Brody et al. (1981); Mehta (1991); Haake (2010); Kos et al. (2018), the Sachdev-Ye-Kitaev model Cotler et al. (2017b, a), and the Hubbard model Zhou et al. (2023); Wittmann W. et al. (2022); Dağ et al. (2023). The intriguing dip-ramp-plateau structure observed in many chaotic systems indicats the universality of SFF Gharibyan et al. (2018); Saad et al. (2019); Winer et al. (2020). The SFF can also be generalized to encapsulate the correlation between multi energy levels Liu (2018); Wei et al. (2024); Cardella (2021); Gross and Rosenhaus (2017).
@@
In quantum systems, the chaotic feature can also be manifested through operator growth. Specifically, the evolution of an initially simple operator $\hat{\cal O}(0)$ gradually becomes more complex under the Heisenberg evolution. This occurs because the operator $\hat{\cal O}(0)$ does not commute with the Hamiltonian $\hat{H}$ of an interacting quantum system. Several approaches have been proposed to quantify the operator growth Nahum et al. (2018); Khemani et al. (2018); Chan et al. (2018b); Roberts et al. (2018). Recently, the concept of K-complexity was introduced to characterize operator growth Parker et al. (2019) and has been applied to various systems, including strongly correlated systems, quantum circuits, and black hole physics Barbón et al. (2020); Magán and Simón (2020); Jian et al. (2021); Kar et al. (2022); Liu et al. (2022); Barbón et al. (2019); Dymarsky and Gorsky (2020); Xu et al. (2020); Avdoshkin and Dymarsky (2020); Rabinovici et al. (2021); Dymarsky and Smolkin (2021); Trigueros and Lin (2022); Kim et al. (2022); Hörnedal et al. (2022); Rabinovici et al. (2022a); Bhattacharjee et al. (2022); Balasubramanian et al. (2022); Heveling et al. (2022); Adhikari and Choudhury (2022); Adhikari et al. (2022); Caputa and Liu (2022); Mück and Yang (2022); Banerjee et al. (2022); Fan (2022a, b); Rabinovici et al. (2022b); Bhattacharyya et al. (2023). It was found that operator growth can be quantified by the mean value of wave packet position on a half-infinite chain. Motivated by this idea, a universal hypothesis of operator growth has also been proposed, where the Lanczos coefficients can be used to diagnose the presence of quantum chaos. Meanwhile, the autocorrelation function ${\cal C}(t)={\rm Tr}(\hat{\cal O}(t)\hat{\cal O}(0))$ defined at infinite temperature can be readily obtained and has been intensively studied in various contexts Kubo (1957); Alhambra et al. (2020); Carabba et al. (2022). It is surprising to find that the universal behavior of the autocorrelation function is determined by the first few Lanczos coefficients Zhang and Zhai (2023). In addition, some studies have explored the relations between K-complexity and other complexity measures, such as circuit complexity, leading to a geometric interpretation of K-complexity Caputa et al. (2022); Lv et al. (2023); Hörnedal et al. (2023).
@@ II Lanczos coefficients at finite temperature
Let us begin with a brief review of the derivation of Lanczos coefficients, and highlight the finite temperature effect of operator growth. Consider a scenario where we have a simple operator $\hat{\cal O}$ and a Hamiltonian $\hat{H}$. Typically, in the presence of interactions, they do not commute. The evolution of the operator $\hat{\cal O}$ follows the Heisenberg equation $-i\partial_{t}\hat{\cal O}(t)=[\hat{H},\hat{\cal O}(t)]$, and the formal solution reads $\hat{\cal O}(t)={\rm e}^{i\hat{H}t}\hat{\cal O}{\rm e}^{-i\hat{H}t}$. Henceforth, we set $\hbar=1$. Therefore, the initial simple operator becomes progressively more complex due to the noncommutative relation between $\hat{\cal O}(t)$ and $\hat{H}$, which is also termed as operator growth. In order to quantify the complexity of $\hat{\cal O}(t)$, the concept of K-complexity was introduced by Parker, et.al. Parker et al. (2019). Given a Hilbert space spanned by |${\ket{i}}$, the operator $\hat{\cal O}=\sum_{i,j}{\cal O}_{ij}\ket{i}\bra{j}$ | can be mapped to an “operator state” $|{\hat{\cal O}}\rangle=\sum_{i,j}{\cal O}_{ij}\ket{i}\otimes\ket{j}$ living in the double Hilbert space spanned by |${\ket{i}\otimes\ket{j}}$. By defining a Liouvillian superoperator ^$\hat{\cal L}=[\hat{H},\cdots]$, the Heisenberg equation can be rewritten as $\partial_{t}|{\hat{\cal O}(t)}\rangle=i\hat{\cal L}|{\hat{\cal O}(t)}\rangle$. In the double Hilbert space, the Liouvillian superoperator appears as a conventional operator, and can be written as ^$\hat{\cal L}=\hat{H}\otimes\hat{\mathbb{I}}-\hat{\mathbb{I}}\otimes\hat{H}^{T}$ where $\hat{\mathbb{I}}$ denotes the identity operator with the same dimension of $\hat{H}$. The superscript $T$ denotes transpose. The time-dependent operator state can be expressed as:
@@
This means that $|{\hat{\cal O}(t)}\rangle$ can be expanded under the basis spanned by |^${|{\hat{\cal L}^{n}\hat{\cal O}}}\rangle$. It is important to note that this basis is neither normalized nor orthogonal. To address this issue, the Krylov basis is defined by the Gram-Schmidt orthogonalization. It is pointed out that one must properly define the inner product of two operator states as follows:
@@
Here, by “standard” we mean that this kind of inner product is widely adopted in the linear response theory. (2) Another inner product often considered in quantum field theory is the Wightman inner product, defined by $g(\lambda)=\delta(\lambda-\beta/2)$, and can be explicitly written as:
@@
It should be noted that at infinitely high temperature, i.e., $\beta\to 0$, both the standard and Wightman inner products reduce to the same one: $\langle{\hat{A}}|{\hat{B}}\rangle={\rm Tr}\left(\hat{A}^{\dagger}\hat{B}\right)/{\rm Tr}(\hat{\mathbb{I}})$. Equipped with the inner product, we can define the Krylov basis denoted by ${|{\hat{\cal O}_{n}}}\rangle$ using the Lanczos algorithm as follows: for $n=1,2$, we have $|{\hat{\cal O}_{0}}\rangle=|{\hat{\cal O}}\rangle$ and $|{\hat{\cal O}_{1}}\rangle=b_{1}^{-1}\hat{\cal L}|{\hat{\cal O}_{0}}\rangle$, respectively, where $b_{1}=\langle{\hat{\cal L}\hat{\cal O}_{0}}|{\hat{\cal L}\hat{\cal O}_{0}}\rangle^{1/2}$; then for $n>2$, $|{\hat{\cal O}_{n}}\rangle$ can be defined inductively as:
@@
Here, $b_{n}$’s indicate the Lanczos coefficients, which obviously depends on temperature. Previous research has demonstrated that the asymptotic behavior of these coefficients at large $n$ can be utilized as a diagnostic for chaotic systems Parker et al. (2019). Using the completeness and orthogonality of the Krylov basis, the Heisenberg equation can be mapped to a discrete Schrödinger-like equation $\partial_{t}\varphi_{n}(t)=b_{n}\varphi_{n-1}(t)-b_{n+1}\varphi_{n+1}(t)$ with $\varphi_{n}(t)=i^{-n}\langle{\cal O}_{n}|{\cal O}(t)\rangle$. The natural initial condition reads $\varphi_{n}(0)=\delta_{n0}$.
@@
To illustrate the difference in ${b_{n}}$’s defined by the standard and Wightman inner products, we calculate ${b_{n}}$ for the Gaussian orthogonal ensemble (GOE), which serves as an ideal chaotic model. The Hamiltonian can be expressed as $\hat{H}_{\rm GOE}=\mathcal{E}{\rm GOE}(L,\mu,\sigma)$, where the $\mathcal{E}$ is the characteristic energy and ${\rm GOE}(L,\mu,\sigma)$ represents a $L\times L$ Hermitian matrix of which the elements are sampled from Gaussian distribution with mean value $\mu$ and standard deviation $\sigma$. The operator is chosen as $\hat{\sigma}_{z}\otimes\hat{\mathbb{I}}_{L/2}$, where $\hat{\sigma}_{z}$ and $\hat{\mathbb{I}}_{L/2}$ denote the Pauli matrix and $L/2\times L/2$ identity matrix, respectively. Obviously, this is a simple operator. By simple, we mean it is represented by a sparse matrix. In evolution, this operator becomes complex progressively. The ${b_{n}}$’s obtained using the Wightman and standard inner products are shown in Fig. 1(a) and (b), respectively. It is important to note that at infinite temperature, both inner products yield the same ${b_{n}}$ values, aligning with our expectations. However, the $b_{n}$’s at finite temperature showcase quite different behaviors for the above-mentioned inner products. The Wightman inner product leads to a linear $b_{n}$ at small $n$, i.e., $b_{n}\approx\alpha(T)n$. The slope $\alpha$ increases as the temperature rises and approaches infinity at infinite temperature. This is reasonable in the sense that at higher temperatures, the speed of operator growth is larger. The range of $n$ that supports linear $b_{n}$ also extends at low temperatures. This indicates that the universal hypothesis is applicable even at finite temperature when the Wightman inner product is adopted. The saturation at large $n$ is due to finite-size effects. In contrast, $b_{n}$ shows oscillatory behavior at finite temperature when the standard inner product is used, and the oscillatory behavior becomes more pronounced at lower temperatures, which is not aligned with the universal hypothesis. In the following analysis, we will focus on the Wightman inner product. In Fig. 1(c), we present the slope $\alpha(T)$ for various temperatures. It is clear that for GOE,
@@
which, on one hand, is consistent with the universal hypothesis on the upper bound of the Lanczos coefficient growth rate $\alpha(T)\leq\pi k_{B}T$, and on the other hand, manifests GOE being the most chaotic model. The slight derivation at high temperature is due to the finite size effect.
@@ III Spectrum form factor
It is evident that as long as the energy difference between two states is $E$, they contribute to ${\cal C}(E)$, regardless of whether they are the nearest neighbors or not. The SFF can be obtained by a Fourier transform as follows:
@@
It has been demonstrated that the “dip-ramp-plateau” structure of SFF is a universal behavior for quantum chaotic systems. In the long time limit, $K(t)\to 1$ according to our definition. Moreover, the SFF can also be generalized to finite temperature by the analytical extension $t\to t-i\beta$, where $\beta=1/(k_{B}T)$ Cotler et al. (2017b). The finite temperature effect of SFF has been investigated in paradigmatic quantum chaotic systems, and the “dip-ramp-plateau” structure survives, which is reproduced in Fig. 2.
Two time scales can be extracted from the SFF Edwards and Thouless (1972); Thouless (1974); Sierant et al. (2020); Šuntajs et al. (2020); one is the Heisenberg time $t_{\rm H}$; the other one is the Thouless time $t_{\rm Th}$. The former one is defined as $t_{\rm H}=2\pi/\Delta$ with $\Delta$ being the mean energy level spacing which exponentially decreases as the system size increases Sierant et al. (2020); Šuntajs et al. (2020). In SFF, the Heisenberg time can be pinned down by the presence of plateau. The Thouless time is the timescale that characterizes how initial local information diffuses throughout the system, and can be defined as the timescale beyond which the SFF becomes consistent the universal one of GOE. In contrast to the Heisenberg time, the Thouless time increases polynomially as the system time increases Sierant et al. (2020); Šuntajs et al. (2020); D’Alessio et al. (2016). In RMT, the analytical form of $K(t)$ at infinite temperature is available. For instance, $K(t)$ of GOE at intermediate and long time can be written as Vasilyev et al. (2020)
$K_{\rm a}(t)=\left\{\begin{aligned} &\frac{2t}{t_{\rm H}}-\frac{t}{t_{\rm H}}{\rm log}_{10}\left(1+\frac{2t}{t_{\rm H}}\right),0<t<t_{\rm H}\\ &2-\frac{t}{t_{\rm H}}{\rm log}_{10}\left(\frac{2t+t_{\rm H}}{2t-t_{\rm H}}\right),t>t_{\rm H}\end{aligned}\right.,$ (9)
where we have taken the random average with respect to the probability of GOE. It can be checked that in the long time limit $K_{\rm a}(t)\to 1$, which is consistent with our definition in Eq.(8).
Figure 2: The spectrum form factor(SFF) at finite temperature. The “dip-ramp-plateau” structure of SFF survives at finite temperature(colored line). The Thouless time(colored diamond) is determined as SFF at finite temperature is close enough to the analytical SFF(black dash line) of GOE at infinite temperature. The Heisenberg time is shown as the black triangle marker.
@@
Similar to the generalization of SFF, we extend the ergodicity indicator to finite temperature since the Heisenberg and Thouless time can still be extracted from the SFF at finite temperature. In Fig. 2, we illustrate the extraction of Heisenberg and Thouless time from SFF at finite temperature. Specifically, we numerically generate 150 samples of GOE at system size of $L=4096$ for each considered temperature. Following the method in Sierant et al. (2020); Šuntajs et al. (2020), we apply the unfolding procedure, which maps each eigenvalue $E_{i}$ to a dimensionless number $\epsilon_{i}$ times average level spacing $\Delta$ and avoids the global energy dependence. Meanwhile, the information in local energy fluctuation reserves Haake (2010). Then the SFF after the unfolding procedure is written as

### v2

@@ Abstract
Unveiling the scaling relations of physical quantities is an important and fundamental topic in the study of classical and quantum physics Stanley (1999); Huckestein (1995); Osterloh et al. (2002); García-García and Wang (2008); Serra et al. (1998); Pal and Huse (2010); Fisher and Berker (1982). In the context of equilibrium systems, significant progress has been made in this direction, leading to a deeper understanding of critical phenomena, phase transitions, and universal behaviors Hung et al. (2011); Puebla et al. (2019); Chen et al. (2019); Chauhan et al. (2022). The investigations of scaling relations in numerous systems are still attracting growing interests, especially in the context of non-equilibrium systems and quantum dynamics of chaotic systems Sierant et al. (2020); Šuntajs et al. (2020).
@@
Typically, the characteristics of quantum chaos manifest across the entire spectrum. To quantify the chaos strength, the statistical properties are conventionally studied. For example, the Wigner-Dyson distribution describes the statistical properties of a quantum chaotic system at infinite temperature, where a level repulsion presents in the energy spectrum Guhr et al. (1998); Mehta (1991); Haake (2010). The statistical properties of the spectrum provide information solely about the nearest level spacing within the spectrum. In contrast, the SFF measures energy correlations between any two states, which allows it to go beyond the constraints of spectrum statistics and capture richer information. SFF has found widespread use in diverse contexts, including random matrix theory (RMT) Guhr et al. (1998); Brody et al. (1981); Mehta (1991); Haake (2010); Kos et al. (2018), the Sachdev-Ye-Kitaev model Cotler et al. (2017b, a), and the Hubbard model Zhou et al. (2024); Wittmann W. et al. (2022); Dağ et al. (2023). The intriguing dip-ramp-plateau structure observed in many chaotic systems indicats the universality of SFF Gharibyan et al. (2018); Saad et al. (2019); Winer et al. (2020). The SFF has also been generalized to encapsulate the correlation between multi energy levels Liu (2018); Wei et al. (2024); Cardella (2021); Gross and Rosenhaus (2017).
@@
In quantum systems, the chaotic feature is also manifested through operator growth. Specifically, the evolution of an initially simple operator $\hat{\cal O}(0)$ gradually becomes more complex under the Heisenberg evolution. This is because the operator $\hat{\cal O}(0)$ does not commute with the Hamiltonian $\hat{H}$ of an interacting quantum system. Several approaches have been proposed to quantify the operator growth Nahum et al. (2018); Khemani et al. (2018); Chan et al. (2018b); Roberts et al. (2018). Recently, the concept of K-complexity was introduced to characterize operator growth Parker et al. (2019) and has been applied to various systems, including strongly correlated systems, quantum circuits, and black hole physics Barbón et al. (2020); Magán and Simón (2020); Kar et al. (2022); Liu et al. (2023); Bhattacharya et al. (2022); Bhattacharjee et al. (2023); Bhattacharya et al. (2023); Bhattacharjee et al. (2024); Barbón et al. (2019); Dymarsky and Gorsky (2020); Xu et al. (2020); Avdoshkin and Dymarsky (2020); Rabinovici et al. (2021); Trigueros and Lin (2022); Kim et al. (2022); Hörnedal et al. (2022); Rabinovici et al. (2022a); Bhattacharjee et al. (2022); Heveling et al. (2022); Adhikari and Choudhury (2022); Adhikari et al. (2023); Caputa and Liu (2022); Mück and Yang (2022); Banerjee et al. (2022); Fan (2022a, b); Rabinovici et al. (2022b); Bhattacharyya et al. (2023); Nandy et al. (2024); Jian et al. (2021); Dymarsky and Smolkin (2021); Balasubramanian et al. (2022). It was found that operator growth can be quantified by the mean value of wave packet position on a half-infinite chain. Motivated by this idea, a universal hypothesis of operator growth has also been proposed, where the Lanczos coefficients can be used to diagnose the presence of quantum chaos. Meanwhile, the autocorrelation function ${\cal C}(t)={\rm Tr}(\hat{\cal O}(t)\hat{\cal O}(0))$ defined at infinite temperature can be readily obtained and has been intensively studied in various contexts Carabba et al. (2022); Kubo (1957); Alhambra et al. (2020). It is surprising to find that the universal behavior of the autocorrelation function is determined by the first few Lanczos coefficients Zhang and Zhai (2024). In addition, some studies have explored the relations between K-complexity and other complexity measures, such as circuit complexity, leading to a geometric interpretation of K-complexity Caputa et al. (2022); Lv et al. (2023); Hörnedal et al. (2023). Furthermore, the concept of K-complexity is extended to systems at finite temperatures, and relevant discussions are presented Parker et al. (2019); Jian et al. (2021); Dymarsky and Smolkin (2021); Balasubramanian et al. (2022); Carabba et al. (2022); Camargo et al. (2023, 2024); Tang (2023).
@@ II Lanczos coefficients at finite temperature
Let us begin with a brief review of the derivation of Lanczos coefficients, and highlight the finite temperature effect of operator growth. Consider a scenario where we have a simple operator $\hat{\cal O}$ and a Hamiltonian $\hat{H}$. Typically, in the presence of interactions, they do not commute. The evolution of the operator $\hat{\cal O}$ follows the Heisenberg equation $-i\partial_{t}\hat{\cal O}(t)=[\hat{H},\hat{\cal O}(t)]$, and the formal solution reads $\hat{\cal O}(t)={\rm e}^{i\hat{H}t}\hat{\cal O}{\rm e}^{-i\hat{H}t}$. Henceforth, we set $\hbar=1$. Therefore, the initial simple operator becomes progressively more complex due to the noncommutative relation between $\hat{\cal O}(t)$ and $\hat{H}$, which is also termed as operator growth. In order to quantify the complexity of $\hat{\cal O}(t)$, the concept of K-complexity was introduced by Parker, et.al. Parker et al. (2019). Given a Hilbert space spanned by |${\ket{i}}$, the operator $\hat{\cal O}=\sum_{i,j}{\cal O}_{ij}\ket{i}\bra{j}$ | can be mapped to an “operator state” $|{\hat{\cal O}}\rangle=\sum_{i,j}{\cal O}_{ij}\ket{i}\otimes\ket{j}$ living in the double Hilbert space spanned by |${\ket{i}\otimes\ket{j}}$. By defining a Liouvillian superoperator ^$\hat{\cal L}=[\hat{H},\cdots]$, the Heisenberg equation can be rewritten as $\partial_{t}|{\hat{\cal O}(t)}\rangle=i\hat{\cal L}|{\hat{\cal O}(t)}\rangle$. In the double Hilbert space, the Liouvillian superoperator appears as a conventional operator, and is written as ^$\hat{\cal L}=\hat{H}\otimes\hat{\mathbb{I}}-\hat{\mathbb{I}}\otimes\hat{H}^{T}$ where $\hat{\mathbb{I}}$ denotes the identity operator with the same dimension of $\hat{H}$. The superscript $T$ denotes transpose. The time-dependent operator state is expressed as:
@@
This means that $|{\hat{\cal O}(t)}\rangle$ is expanded under the basis spanned by |^${|{\hat{\cal L}^{n}\hat{\cal O}}}\rangle$. It is important to note that this basis is neither normalized nor orthogonal. To address this issue, the Krylov basis is defined by the Gram-Schmidt orthogonalization. It is pointed out that one must properly define the inner product of two operator states as follows:
@@
Here, by “standard” we mean that this kind of inner product is widely adopted in the linear response theory. (2) Another inner product often considered in quantum field theory is the Wightman inner product, defined by $g(\lambda)=\delta(\lambda-\beta/2)$, and is explicitly written as,
@@
It should be noted that at infinitely high temperature, i.e., $\beta\to 0$, both the standard and Wightman inner products reduce to the same one: $\langle{\hat{A}}|{\hat{B}}\rangle={\rm Tr}\left(\hat{A}^{\dagger}\hat{B}\right)/{\rm Tr}(\hat{\mathbb{I}})$. With the inner product, we define the Krylov basis ${|{\hat{\cal O}_{n}}}\rangle$ using the Lanczos algorithm as follows: for $n=1,2$, we have $|{\hat{\cal O}_{0}}\rangle=|{\hat{\cal O}}\rangle$ and $|{\hat{\cal O}_{1}}\rangle=b_{1}^{-1}\hat{\cal L}|{\hat{\cal O}_{0}}\rangle$, respectively, where $b_{1}=\langle{\hat{\cal L}\hat{\cal O}_{0}}|{\hat{\cal L}\hat{\cal O}_{0}}\rangle^{1/2}$; then for $n>2$, $|{\hat{\cal O}_{n}}\rangle$ is defined inductively as,
@@
Here, $b_{n}$’s indicate the Lanczos coefficients, which obviously depend on temperature. Previous research has demonstrated that the asymptotic behavior of these coefficients at large $n$ serves as a diagnostic for chaotic systems Parker et al. (2019). Using the completeness and orthogonality of the Krylov basis, the Heisenberg equation maps to a discrete Schrödinger-like equation $\partial_{t}\varphi_{n}(t)=b_{n}\varphi_{n-1}(t)-b_{n+1}\varphi_{n+1}(t)$ with $\varphi_{n}(t)=i^{-n}\langle{\cal O}_{n}|{\cal O}(t)\rangle$. The natural initial condition reads $\varphi_{n}(0)=\delta_{n0}$.
@@
To illustrate the difference in ${b_{n}}$’s defined by the standard and Wightman inner products, we calculate ${b_{n}}$ for the Gaussian orthogonal ensemble (GOE), which is an ideal chaotic model. The Hamiltonian is expressed as $\hat{H}_{\rm GOE}={\rm GOE}(L,\mu,\sigma)$, where ${\rm GOE}(L,\mu,\sigma)$ represents a $L\times L$ Hermitian matrix of which the elements are sampled from Gaussian distribution with mean value $\mu$ and standard deviation $\sigma$. In our case, we choose $\mu=0$ and $\sigma=\mathcal{E}$, where the $\mathcal{E}$ is the characteristic energy. The operator is chosen as $\hat{\sigma}_{z}\otimes\hat{\mathbb{I}}_{L/2}$, where $\hat{\sigma}_{z}$ and $\hat{\mathbb{I}}_{L/2}$ denote the Pauli matrix and $L/2\times L/2$ identity matrix, respectively. Obviously, this is a simple operator. By simple, we mean it is represented by a sparse matrix. In evolution, this operator becomes complex progressively. The ${b_{n}}$’s obtained using the Wightman and standard inner products are shown in Fig. 1(a) and (b), respectively. It is important to note that at infinite temperature, both inner products yield the same ${b_{n}}$ values, aligning with our expectations. However, the $b_{n}$’s at finite temperature showcase quite different behaviors for the above-mentioned inner products. The Wightman inner product leads to a linear $b_{n}$ at small $n$, i.e., $b_{n}\approx\alpha(T)n$. The slope $\alpha$ increases as the temperature rises and approaches infinity at infinite temperature. This is reasonable in the sense that at higher temperatures, the speed of operator growth is larger. The range of $n$ that supports linear $b_{n}$ also extends at low temperatures. This indicates that the universal hypothesis is applicable even at finite temperature when the Wightman inner product is adopted. The saturation at large $n$ is due to finite-size effects. In contrast, $b_{n}$ shows oscillatory behavior at finite temperature when the standard inner product is used, and the oscillatory behavior becomes more pronounced at lower temperatures, which is not aligned with the universal hypothesis. In the following analysis, we will focus on the Wightman inner product. In Fig. 1(c), we present the slope $\alpha(T)$ for various temperatures. It is clear that for GOE,
@@
which, on the one hand, is consistent with the universal hypothesis on the upper bound of the Lanczos coefficient growth rate $\alpha(T)\leq\pi k_{B}T$, and on the other hand, manifests GOE being the most chaotic model. The small derivation at high temperature is due to the finite size effect.
@@ III Spectrum form factor
It is evident that as long as the energy difference between two states is $E$, they contribute to ${\cal C}(E)$, regardless of whether they are the nearest neighbors or not. The SFF is obtained by a Fourier transformation as follows,
@@
It has been demonstrated that the “dip-ramp-plateau” structure of SFF is a universal behavior for quantum chaotic systems. In the long time limit, $K(t)\to 1$ according to our definition. Moreover, the SFF has also been generalized to finite temperature by the analytical extension $t\to t-i\beta$, where $\beta=1/(k_{B}T)$ Cotler et al. (2017b). The finite temperature effect of SFF has been investigated in paradigmatic quantum chaotic systems, and the “dip-ramp-plateau” structure survives, which is reproduced in Fig. 2.
The SFF is characterized by two time scales  Edwards and Thouless (1972); Thouless (1974); Sierant et al. (2020); Šuntajs et al. (2020); one is the Heisenberg time $t_{\rm H}$; the other one is the Thouless time $t_{\rm Th}$. The former one is defined as $t_{\rm H}=2\pi/\Delta$ with $\Delta$ being the mean energy level spacing which exponentially decreases as the system size increases Sierant et al. (2020); Šuntajs et al. (2020). In SFF, the Heisenberg time can be pinned down by the presence of plateau. The Thouless time is the timescale that characterizes how initial local information diffuses throughout the system, and is defined as the timescale beyond which the SFF becomes consistent the universal one of GOE. In contrast to the Heisenberg time, the Thouless time increases polynomially as the system size increases Sierant et al. (2020); Šuntajs et al. (2020); D’Alessio et al. (2016). In RMT, the analytical form of $K(t)$ at infinite temperature is available. For instance, $K(t)$ of GOE at intermediate and long time is written as Vasilyev et al. (2020)
$K_{\rm a}(t)=\left\{\begin{aligned} &\frac{2t}{t_{\rm H}}-\frac{t}{t_{\rm H}}{\rm ln}\left(1+\frac{2t}{t_{\rm H}}\right),0<t<t_{\rm H}\\ &2-\frac{t}{t_{\rm H}}{\rm ln}\left(\frac{2t+t_{\rm H}}{2t-t_{\rm H}}\right),t>t_{\rm H}\end{aligned}\right.,$ (9)
where we have taken the random average with respect to the probability of GOE. In the long time limit $K_{\rm a}(t)\to 1$, which is consistent with our definition in Eq.(8).
Figure 2: The spectrum form factor(SFF) at finite temperature. The “dip-ramp-plateau” structure of SFF survives at finite temperature(colored line). The Thouless time(colored diamond) is determined as SFF at finite temperature is close enough to the analytical SFF(black dash line) of GOE at infinite temperature. It should be noticed that, by definition, the Thouless time is different from the dip time when the SFF reaches its minimum value. The Heisenberg time is shown as the black triangle marker.
@@
Similar to the generalization of SFF, we extend the ergodicity indicator to finite temperature since the Heisenberg and Thouless time can still be extracted from the SFF at finite temperature. In Fig. 2, we illustrate the extraction of Heisenberg and Thouless time from SFF at finite temperature. Specifically, we numerically generate 150 samples of GOE at system size of $L=4096$ for each considered temperature. Following the method in Sierant et al. (2020); Šuntajs et al. (2020), we apply the unfolding procedure, which maps each eigenvalue $E_{i}$ to a dimensionless number $\epsilon_{i}$ times average level spacing $\Delta$ and avoids the global energy dependence. Meanwhile, the information in local energy fluctuations is preserved Haake (2010). Then the SFF after the unfolding procedure is written as